OpenStudy (anonymous):

A cruise ship left Diego Garcia at the same time as the container ship. The ships traveled in opposite directions. The container ship traveled at a speed of 25 km/h. After 13 hours they were 650 km apart. How fast did the cruise ship travel?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Thats what I thought '=, THANKS!

hero (hero):

13*25 = 325 650 - 325 = 325 325/13 = 25 So both ships travelled at the same speed.